Adding drugs to IV fluids A 10-kg dog has been prescribed a drug with the strength of 6 mg/ml infuse the dog is 2 mgkg/day (24 hours). The dog is receiving maintenance fluids at a maintenancc rate of 50 ml/kg/day. The drug is to be added to a l-L bag of tluid a. Calculate the amount of drug to add to the fluid. What is the drip rate over which the Auid should be given using a giving set that
Answer #1

1. Weight of the dog is 10kg

Strength of available drug is 6mg/ml

The rate to infuse the dog is 2mg/kg/day

2mg/kg= 2×10 =20mg of drug has to be infused to the dog a day.

Now, 6mg of drug is present in 1ml

1mg of drug will be present in 1/6ml

Then 20mg of drug will be present in 1/6 × 20 = 3.33ml

So, 3.33ml of drug has to be added into the fluid .

2. Drip rate = total volume /total time in mins ×drop factor

Drop factor =60

The dog is receiving maintenance fluid at 50mg/kg/day

So the total volume to be infused 50×20 ml a day.

= 1000ml

Total time =24hr = 24×60 mins

Drip rate = 1000 ×60

24×60

= 1000/24 ml/min

= 41.6 ml/min is the drip rate